OpenStudy (anonymous):

simplify (x^3/5)^2/3 Answer choices: A.x^9/125 B.x^5/8 C.x^2/5 D.x^6/15 please help i give medals!

OpenStudy (anonymous):

\[(\frac{ x^6 }{ 5 })^{2/3}\] Is this is?

OpenStudy (anonymous):

No its (x^3/5)^2/3

OpenStudy (anonymous):

oops, wrong number. But yea all you do is multiply the exponent out

OpenStudy (anonymous):

do you know which one it would be?

OpenStudy (anonymous):

i thought c

OpenStudy (anonymous):

Well c seems like it's correct

OpenStudy (anonymous):

although the denominator seems to be the problem

OpenStudy (goldphenoix):

\[\large \large \frac{ (x^3/5)^2 }{ 3 }\] ^Does it look like this?

OpenStudy (anonymous):

(x^3/5)^2/3) exactley like that

OpenStudy (anonymous):

exactly*

OpenStudy (anonymous):

Then I don't think any of the answer choices match

OpenStudy (anonymous):

the 3/5 is raised like an exponet of the X and the 2/3 is an exponet of the whole thing

OpenStudy (goldphenoix):

\[\large \large \frac{ (x^3/5)(x^3/5) }{ 3 } \]

OpenStudy (goldphenoix):

You know how to solve this?

OpenStudy (jhannybean):

It helps when you use brackets to differentiate where your numerator and denominator are.

OpenStudy (anonymous):

thanks ! So its a?

OpenStudy (anonymous):

I don't think 25*6=125

OpenStudy (anonymous):

how do yall write in the fancy math text, ill put it into that

OpenStudy (goldphenoix):

You click the equation button.

OpenStudy (jhannybean):

\[\large \left (\frac{x^3}{5}\right )^2\cdot \left ( \frac{1}{3}\right )\]\[\large \left ( \frac{x^{3\cdot 2}}{5^2} \right ) \cdot \frac13 \]\[\large \left (\frac{x^6}{25}\right ) \cdot \frac13\]\[\large \frac{x^6}{25 \cdot 3} = ?\]

OpenStudy (goldphenoix):

Jhanny is correct.

OpenStudy (anonymous):

are you sure your answer choices are correct

OpenStudy (jhannybean):

your answer choices are wrong, i believe.

OpenStudy (goldphenoix):

Yes. Jhanny is 100 percent correct. The answer choice is wrong.

OpenStudy (jhannybean):

when you square a fraction, you distribute the square to the numerator and denominator

OpenStudy (anonymous):

i dont know how to write the equation write on here But the equation you wrote wasnt in the write form

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):